# How do Laws Produce the Future?

### v1

@@ 3 Newtonian Gravity
One might object that Newtonian gravity is supposed to be a time-symmetric theory, but the earlier characterization of Newton’s second law in terms of future acceleration (3) is time-asymmetric. That law tells you how velocities change towards the future, not the past. Let me make three remarks about this time-asymmetry. First, although standard presentations of the equations of Newtonian gravity may not display any such time-asymmetry, explanations of those equations are generally time-asymmetric. Forces are described as causing future acceleration, an understanding that fits well with (3) (Easwaran,, 2014, pg. 852–853). Second, the theory remains time-symmetric in the sense that it is time-reversal invariant: for any history that obeys (3) and (4), the time-reversed history will obey (3) and (4). The key point here is that any history where $\vec{F}=m\vec{a}^{\,pf}$ is also a history where $\vec{F}=m\vec{a}^{\,fp}$ because $\vec{a}^{\,pf}$ cannot differ from $\vec{a}^{\,fp}$ if particles are only interacting through well-behaved gravitational forces. Third, given that we can introduce the above time-directed version of the Newton’s second law (3) and view the past as generating the future, there is still the question as to whether we should. There we enter the broader debate between the dynamic production account and its rivals.
The above story about the laws of Newtonian gravity changes if you take the ontology (what exists according to the theory) to include a gravitational field $\vec{g}$ in addition to the point masses. Then, you might try the following collection of laws:
@@
The first law is Newton’s second law, appearing again as the only dynamical law. The second law relates the force on a body to the gravitational field at the body’s location $\vec{x}_{i}$. The third law gives the gravitational field resulting from a certain arrangement of masses as a sum of contributions from each mass. It is interesting to note that the gravitational field does not have its evolution determined by a dynamical law. It is rebuilt at each moment by a non-dynamical law.
The set of laws for Newtonian gravity (including a gravitational field) that were just given (6) will not function properly because they have problems with self-interaction. As you approach the location of any point mass, the gravitational field becomes infinitely strong and does not have a clear direction (figure 3). The field is ill-defined at the mass’s location, which is exactly where you would like to use it to calculate the force on the mass. One way to address this problem is to introduce a separate gravitational field sourced by each particle and stipulate that particles do not feel their own fields. Then, you could derive Newton’s law of gravitation (4) with the important $j\neq i$ in the sum. A different strategy for addressing self-interaction would be to retain a single gravitational field and replace point masses by extended bodies, either rigid bodies or continua. [^16] For discussion of the different possible ontologies for classical mechanics (point particles, rigid bodies, and continua), see Wilson, (1998, 2013); van Strien, (2021). Then, we would have laws like:
@@
where $G$ is the gravitational constant, $\rho^{m}$ is the mass density, and $\vec{f}$ is the density of force exerted by the gravitational field on matter. (Note that because Newtonian gravity closely parallels electrostatics, the first two laws resemble Maxwell’s equations as they apply to electrostatics.) None of the three laws presented above are dynamical laws. To get dynamics, we would also need a version of Newton’s second law. For rigid bodies, one can integrate the force density over the volume of the body to get a net force and use this to determine how the center of mass moves. One would also need a law connecting the density of force on the body to its rotation. For continua, the reaction to experienced forces will depend on the nature of the body experiencing the forces. (We will return to that issue in section 6.) Eliminating the gravitational field, you could alternatively have rigid bodies or continua interacting by gravitational action-at-a-distance—as in (4). All of this begins to illustrate the plethora of options that you have for understanding the laws and ontology of Newtonian gravity while staying within the time evolution paradigm. [^17] There are also ways of formulating the laws that would fall outside the time evolution paradigm, as in the Lagrangian stationary action approach (Wharton,, 2015; Adlam,, 2022a, pg. 3–4, 35; Chen & Goldstein,, 2022, pg. 46–47).
@@ 4 Electromagnetism, First Pass
Although the retarded action-at-a-distance and Wheeler-Feynman versions of electromagnetism are worthy of study, there are good reasons to prefer a version of electromagnetism that takes the electromagnetic field to be real. Here are three: First, a real electromagnetic field can ensure conservation of energy and momentum by itself possessing energy and momentum. [^19] Lazarovici, (2018, sec. 4.2) gives a response to this concern defending the Wheeler-Feynman theory. Lange, (2002) presents the argument for the reality of the electromagnetic field from conservation of energy and momentum, but finds it lacking in a relativistic context because energy and momentum are frame-dependent quantities. He bases his argument for the field’s reality on the fact that it possesses mass. (Physicists classify the electromagnetic field as a massless field because the associated particle—the photon—is massless, but there is an important sense in which the field does carry mass in addition to energy and momentum. See Lange,, 2002, ch. 8; Sebens,, 2022b, sec. 2.) Second, in general relativity the electromagnetic field acts as a source for gravitation. Third, in quantum electrodynamics the electromagnetic field is treated in a similar way to charged matter and thus it is natural (with the benefit of hindsight) to view both as real in classical electromagnetism (Sebens,, 2022a; Hubert & Sebens,, 2023). [^20] Proponents of the Wheeler-Feynman theory can of course pursue a revisionary approach to quantum field theory that eliminates the electromagnetic field and thus does not treat it in a similar way to charged matter.
@@
Jefimenko takes Maxwell’s equations to express relations between present-time quantities. (We will see shortly that, by interpreting the time derivatives as future derivatives, two of Maxwell’s equation can actually be understood as linking cause and effect with cause preceding effect. [^21] This response to Jefimenko is similar to the response that Easwaran, (2014, pg. 853) gives to Field,’s (2003, pg. 438) general claim that the differential equations that appear in laws of physics cannot be interpreted causally because they relate quantities at the same time.) Unlike Maxwell’s equations, Jefimenko’s equations connect present fields to past sources and thus straightforwardly fit his mold. However, that connection is temporally non-local and thus, although they may conform to our basic expectation that causes should come before effects, Jefimenko’s equations do not fit the time evolution paradigm as stated in section 2.
The restricted version of electromagnetism that Jefimenko prefers forbids unsourced radiation. It thus appears to yield a natural explanation as to why electromagnetic waves diverge—an explanation of the arrow of electromagnetic radiation. That asymmetry can also be explained in an unrestricted version of electromagnetism, as it is improbable that any unsourced contribution to the electromagnetic field would result in coordinated converging waves (North,, 2003; Hubert & Sebens,, 2023). There is room to debate which version of electromagnetism gives a better explanation of this asymmetry, but I prefer the explanation offered by an unrestricted version because the same kind of explanation can be given as to why entropy increases and why other waves diverge. An unrestricted version of electromagnetism also fits well with quantum field theory, where the quantum electromagnetic field is not restricted to always have past sources (where, put another way, photons no more need past sources than electrons do). [^22] See Wald, (2022, sec. 1.3); Hubert & Sebens, (2023, sec. 3.3).
@@
Note that here we are, for the moment, adopting an ontology of point particles for the charged matter and thus the charge and current densities in (10) and (13) will feature delta functions (as in Landau & Lifshitz,, 1971, sec. 28). The use of point particles will lead to trouble with self-interaction because the electromagnetic field will become infinitely strong at the location of any particle (just where the field’s value matters for calculating forces). I will set these problems aside for now and we will return to them in section 6. To remove infinities, it would be better to work with charge distributions and avoid point charges.
@@ 5 The Relativistic Time Evolution Paradigm
Although it is not difficult to find a version of electromagnetism that fits the time evolution paradigm, that paradigm may be too loose. Electromagnetism is a relativistic theory and as such we can put tighter constraints on the form that the dynamical laws should take. They should be spatiotemporally local, [^24] Here I am talking about spatiotemporal locality in the special relativistic sense, not the broader sense in which there is no light-speed limit on influences and the only constraint is that there be spatiotemporal continuity (see Lange,, 2002, ch. 1). not merely temporally local. Here is how Maudlin, (2007, pg. 60) puts the requirement of relativistic locality: “the physical state at any point of space-time is determined or influenced only by events in its past light-cone.” [^25] See also Maudlin, (2007, pg. 20 & 30). If the state at any point can only be influenced by events in its past light-cone, then the state at a particular space-time point can only influence events in the future light-cone of that point (for these are the events that it is in the past light-cone of). Given Bell’s theorem, quantum physics may force us to give up on this kind of locality (a point that we will return to in section 7). But, it seems like this kind of locality should at least be achievable within classical electromagnetism.
@@
    Spatiotemporally Local Dynamical Laws: A subset of the laws (the dynamical laws) can be applied at any point in space-time, giving past-to-future light-cone neighborhood properties as output and taking as input only space-time point properties or past light-cone neighborhood properties. The dynamical laws may give either precise values for the past-to-future light-cone neighborhood properties or probability distributions over such values.
@@
    Deterministic or Stochastic: Once a law-abiding history over an arbitrarily small time interval to the past of a given moment has been fixed (in some inertial reference frame), the laws either uniquely fix a single future sequence of states or the laws yield a precise probability distribution over future sequences of states.
Present properties (that make no reference to other times) have been replaced by space-time point properties (that make no reference to other times or places). Past neighborhood properties have been replaced by past light-cone neighborhood properties (that are defined in terms of arbitrarily small subregions of the past light-cone). Similarly, past-to-future properties have been replaced by past-to-future light-cone properties. The third condition acknowledges that the selection of a thin time slice will depend on the frame of reference and requires deterministic or stochastic dynamics going forward from any such slice. The three conditions are not meant to specify what it takes for a theory to be relativistic or to impose the strictest standards on relativistic laws. The conditions are just meant to give a baseline as to the way that you might expect time evolution to work in a relativistic theory that includes laws of time evolution. In particular, the third condition could be strengthened to demand a more local form of determinism or stochasticity. In a relativistic theory, specifying what is going on at a time within any spherical region (and its arbitrarily-short past light-cone) should either fix, or give probabilities for, future slices of the converging light-cone that the spherical region forms the base of. [^26] The source-free homogeneous wave equation achieves this kind of determinism, as is proved in Strauss, (2008, sec. 9.1) and will be discussed in section 6.
@@ 6 Electromagnetism, Second Pass
It requires a bit of work to fit electromagnetism into the relativistic time evolution paradigm, but the trick can be done. Here I present one way to do so, without claiming that it is the best and final way. We will need to find suitable laws governing the evolution of the electromagnetic field, the forces on charged matter, and the reaction of matter to forces. Let us begin with the field. We can use the scalar potential $\phi$ and the vector potential $\vec{A}$ to specify the state of the electromagnetic field, where these potentials are related to $\vec{E}$ and $\vec{B}$ by
@@
Although this first attempt fails, we should be optimistic that the wave equations in (17) and (18) can be interpreted as giving spatiotemporally local dynamics. When the homogenous (source-free) wave equation, $\Box u=0$, is discussed in textbooks on partial differential equations, it is standard practice to prove a “causality theorem” showing that a given spacetime point cannot influence anything outside its future light-cone and cannot be influenced by anything outside its past light-cone. [^29] For discussion of causality theorems for one and three-dimensional wave equations, see Zachmanoglou & Thoe, (1976, ch. 8); Folland, (1995, ch. 5); Evans, (1998, sec. 2.4.3); Strauss, (2008, sec. 9.1). For discussion of causality theorems in electromagnetism (making use of the Lorenz gauge), see Wald, (1984, sec. 10.2); Wald, (2022, sec. 5.4). To be more precise, you can prove that if two solutions agree on the values of the function $u$ and its time derivatives $\frac{\partial u}{\partial t}$ at a given time within the spherical spatial region bounded by the past light-cone of some future spacetime point $(\vec{x},t)$, the two solutions will agree on the value of $u$ at $(\vec{x},t)$. Solutions to the inhomogeneous (sourced) wave equation, $\Box u=f$, will also satisfy such a causality theorem for a fixed source $f$ because any solution can be divided into a free part $u_{in}$ (that obeys the homogenous wave equation and can be interpreted as describing incoming waves that are not attributable to the source function $f$) and a retarded part $u_{ret}$ (where the value of $u_{ret}$ at any space-time point is fixed by $f$ along the past light-cone of that point). Given these causality results, it seems like there should be a way to formulate inhomogeneous wave equations, like the wave equations for the potentials (17) and (18), that makes the causal structure manifest and satisfies the conditions that we laid out earlier for spatiotemporally local dynamics linking past light-cone to future light-cone. As we will see, this can be done straightforwardly in one-dimensional space, but is surprisingly difficult in three-dimensional space.
@@
with $f$ treated as a fixed source function defined across space and time. D’Alembert’s method for solving the homogeneous wave equation introduces new variables
@@
(These are one-dimensional versions of Dirac’s light-cone, or light-front, coordinates. [^30] See Dirac, (1949).) We can use these variables to rewrite (23) as
@@
where the derivatives are evaluated along the two perpendicular edges of the light-cone. We can ask whether these derivatives should be evaluated by considering a small amount of the future or the past. If you were to take both derivatives to be future derivatives, then the output (the left-hand side of the equation) would be a future light-cone neighborhood property (and thus a past-to-future light-cone neighborhood property) and the input (the right-hand side) would be a space-time point property (assuming that $f$ gives space-time point properties). However, this interpretation of (25) would not capture the causal structure accurately. The source function $f$ alone does not determine how $u$ will change towards the future. At a moment, you need $f$, $u$, and the rate at which $u$ is changing to fix the future evolution.
@@
Looking back at (26), you could just as well swap the order of the derivatives and go up the other edge of the future light-cone. To arrive at a more symmetric law and avoid privileging either side of the future light-cone, you could average the two orderings:
@@
where we will again treat $f$ as a fixed source function defined across space and time. The wave equations for the potentials, (17) and (18), fit this general form. Writing this wave equation as a spatiotemporally local dynamical law is not as straightforward as you might expect, but here is one way to do it. Let us focus our attention on the evolution at the origin, noting that any point in space can be treated as the origin. Using spherical coordinates with $\theta$ as the azimuthal angle and $\phi$ as the polar angle, the wave equation (30) for $u$ at the origin becomes
@@
where $\int\frac{d\theta d\phi}{4\pi}\sin\phi$, or $\int\frac{d\Omega}{4\pi}$, integrates over the solid angle $\Omega$ and divides by the total solid angle. For the time derivative term, this integral is trivial. To see that the Laplacian in (30) can be replaced by a radial second derivative in (31), plug the expansion
$\frac{\partial}{\partial r}=\cos\theta\sin\phi\frac{\partial}{\partial x}+\sin\theta\sin\phi\frac{\partial}{\partial y}+\cos\phi\frac{\partial}{\partial z}$ (32)
into the integral over the solid angle. The cross terms drop out and you are left with
$\int\frac{d\Omega}{4\pi}\left(3\frac{\partial^{2}}{\partial r^{2}}\right)u$ $=\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}\right)u=\nabla^{2}u\ .$ (33)
@@
In the integral, for a given $\theta$ and $\phi$, the $\frac{\partial^{2}}{\partial\eta^{2}}$ term asks you to take a second derivative along the future light-cone in that direction and the $\frac{\partial^{2}}{\partial\xi^{2}}$ term asks you to take a second derivative along the past light-cone in that direction. This past light-cone derivative should give the same result as taking the forward light-cone derivative in the opposite direction, $-\hat{r}$. Integrating over all angles, [^32] Here is a more formal proof of (37). Expanded in terms of $r$ and $t$ derivatives, the difference between the left and right-hand sides of (37) is proportional to $\int\frac{d\Omega}{4\pi}\left(\frac{\partial}{\partial r}\frac{\partial}{\partial t}\right)=0\ .$ (36) The fact that this term vanishes can be seen by noting that radial derivatives in opposite directions will cancel, or, by using the expansion of $\frac{\partial}{\partial r}$ in (32).
@@
Let us now distinguish past and future derivatives as in (26),
@@
For given angles $\theta$ and $\phi$, the integrand asks to consider how a sum of two past light-cone derivatives (one in the $\hat{r}$ direction and one in the $-\hat{r}$ direction) changes as you move along the future light-cone in the $\hat{r}$ direction (see figure 6). With this form of the wave equation, we have arrived at a spatiotemporally local dynamical law where the left-hand side is a past-to-future light-cone neighborhood property and the source function on the right-hand side is either a space-time point property or a past light-cone property (as we are about to see). [^33] One should not read the speed of causal influence directly from the ability to put the wave equation in a spatiotemporally local form like (40). If you started with a three-dimensional wave equation (30) with the speed of light of light doubled ($c\rightarrow 2c$), then causal influences would propagate in a way that violates a speed-of-light-based standard for spatiotemporal locality. However, you can follow parallel reasoning to the derivation above to get an expression for the wave equation at the origin in terms of the old coordinates $\eta$ and $\xi$ defined in terms of the original speed of light $c$ as $\int\frac{d\Omega}{4\pi}\left(\left(\frac{\partial}{\partial\eta}\right)^{f}\left[\frac{11}{2}\left(\frac{\partial}{\partial\eta}\right)^{p}+\frac{13}{2}\left(\frac{\partial}{\partial\xi}\right)^{p}\right]\right)u$ $=f\ ,$ (41) which meets our earlier standard for being a spatiotemporally local dynamical law. We must reject this formulation as not accurately capturing the way that the past generates the future for the modified wave equation, but perhaps more could be said as to why that is so.

### v2

@@ 3 Newtonian Gravity
One might object that Newtonian gravity is supposed to be a time-symmetric theory, but the earlier characterization of Newton’s second law in terms of future acceleration (3) is time-asymmetric. That law tells you how velocities change towards the future, not the past. A couple remarks on this asymmetry: First, although standard presentations of the equations of Newtonian gravity may not display any such time-asymmetry, explanations of those equations are generally time-asymmetric. Forces are described as causing future acceleration, an understanding that fits well with (3) (Easwaran,, 2014, pg. 852–853). Second, the theory remains time-symmetric in the sense that it is time-reversal invariant: for any history that obeys (3) and (4), the time-reversed history will obey (3) and (4). The key point here is that any history where $\vec{F}=m\vec{a}^{\,pf}$ is also a history where $\vec{F}=m\vec{a}^{\,fp}$ because $\vec{a}^{\,pf}$ cannot differ from $\vec{a}^{\,fp}$ if particles are only interacting through well-behaved gravitational forces.
The above story about the laws of Newtonian gravity changes if you take the ontology (what exists according to the theory) to include a gravitational field $\vec{g}$ in addition to the point masses. To include the field, you might try the following simple (but ultimately unsatisfactory) collection of laws:
@@
The first law is Newton’s second law, as in (3). The second law relates the force on a body to the gravitational field at the body’s location $\vec{x}_{i}$. The third law gives the gravitational field resulting from a certain arrangement of masses as a sum of contributions from each mass. It is interesting to note that the gravitational field does not have its evolution determined by a dynamical law. It is rebuilt at each moment by a non-dynamical law.
The set of laws for Newtonian gravity (including a gravitational field) that were just given (6) will not function properly because they have problems with self-interaction. As you approach the location of any point mass, the gravitational field becomes infinitely strong and does not have a clear direction (figure 3). The field is ill-defined at the mass’s location, which is exactly where you would like to use it to calculate the force on the mass. One way to address this problem is to introduce a separate gravitational field sourced by each particle and stipulate that particles do not feel their own fields. Then, you could derive Newton’s law of gravitation (4) with the important $j\neq i$ in the sum.
A different strategy for addressing self-interaction would be to retain a single gravitational field and replace point masses by extended bodies, either rigid bodies or continua. [^17] For discussion of the different possible ontologies for classical mechanics (point particles, rigid bodies, and continua), see Wilson, (1998, 2013); van Strien, (2021). Then, we would have laws like:
@@
where $G$ is the gravitational constant, $\rho^{m}$ is the mass density, and $\vec{f}$ is the density of force exerted by the gravitational field on matter. (Note that because Newtonian gravity closely parallels electrostatics, the first two laws resemble Maxwell’s equations as they apply to electrostatics.) None of the three laws presented above are dynamical laws. To get dynamics, we would also need a version of Newton’s second law. For rigid bodies, one can integrate the force density over the volume of the body to get a net force and use this to determine how the center of mass moves. One would also need a law connecting the density of force on the body to its rotation. For continua, the reaction to experienced forces will depend on the nature of the body experiencing the forces. (We will return to that issue in section 6.)
Eliminating the gravitational field, you could alternatively have rigid bodies or continua interacting by gravitational action-at-a-distance—as in (4). All of this begins to illustrate the plethora of options that you have for understanding the laws and ontology of Newtonian gravity while staying within the time evolution paradigm. [^18] There are also ways of formulating the laws that would fall outside the time evolution paradigm, as in the Lagrangian stationary action approach (Wharton,, 2015; Adlam,, 2022a, pg. 3–4, 35; Chen & Goldstein,, 2022, pg. 46–47).
@@ 4 Electromagnetism, First Pass
Although the retarded action-at-a-distance and Wheeler-Feynman versions of electromagnetism are worthy of study, there are good reasons to prefer a version of electromagnetism that takes the electromagnetic field to be real. Here are three: First, a real electromagnetic field can ensure conservation of energy and momentum by itself possessing energy and momentum. [^20] Lazarovici, (2018, sec. 4.2) gives a response to this concern defending the Wheeler-Feynman theory., [^21] Lange, (2002) presents the argument for the reality of the electromagnetic field from conservation of energy and momentum, but finds it lacking in a relativistic context because energy and momentum are frame-dependent quantities. He bases his argument for the field’s reality on the fact that it possesses mass. (Physicists classify the electromagnetic field as a massless field because the associated particle—the photon—is massless, but there is an important sense in which the field carries mass in addition to energy and momentum. See Lange,, 2002, ch. 8; Sebens,, 2022b, sec. 2.) Second, in general relativity the electromagnetic field acts as a source for gravitation. Third, in quantum electrodynamics the electromagnetic field is treated in a similar way to charged matter and thus it is natural (with the benefit of hindsight) to view both as real in classical electromagnetism (Sebens,, 2022a; Hubert & Sebens,, 2023). [^22] Proponents of the Wheeler-Feynman theory can of course pursue a revisionary approach to quantum field theory that eliminates the electromagnetic field and thus does not treat it in a similar way to charged matter.
@@
Jefimenko takes Maxwell’s equations to express relations between present-time quantities. (We will see shortly that, by using future derivatives, two of Maxwell’s equations can be understood as linking cause and effect with cause preceding effect. [^23] This response to Jefimenko is similar to the response that Easwaran, (2014, pg. 853) gives to Field,’s (2003, pg. 438) general claim that the differential equations that appear in laws of physics cannot be interpreted causally because they relate quantities at the same time.) Unlike Maxwell’s equations, Jefimenko’s equations connect present fields to past sources and thus straightforwardly fit his mold. However, that connection is temporally non-local and thus, although they may conform to our basic expectation that causes should come before effects, Jefimenko’s equations do not fit the time evolution paradigm as stated in section 2.
The restricted version of electromagnetism that Jefimenko prefers forbids unsourced radiation. It thus appears to yield a natural explanation as to why electromagnetic waves diverge—an explanation of the arrow of electromagnetic radiation. That asymmetry can also be explained in an unrestricted version of electromagnetism, as it is improbable that any unsourced contribution to the electromagnetic field would result in coordinated converging waves (North,, 2003; Hubert & Sebens,, 2023). There is room to debate which version of electromagnetism gives a better explanation of this asymmetry, but the explanation offered by an unrestricted version has the advantage that the same kind of explanation can be given as to why entropy increases and why other waves diverge. An unrestricted version of electromagnetism also fits well with quantum field theory, where the quantum electromagnetic field is not restricted to always have past sources (where, put another way, photons no more need past sources than electrons do). [^24] See Wald, (2022, sec. 1.3); Hubert & Sebens, (2023, sec. 3.3).
@@
Note that here we are, for the moment, adopting an ontology of point particles for the charged matter and thus the charge and current densities in (10) and (13) will feature delta functions (as in Landau & Lifshitz,, 1971, sec. 28). The use of point particles will lead to trouble with self-interaction, as in section 3, because the electromagnetic field will become infinitely strong at the location of any particle (just where the field’s value matters for calculating forces). Let us set these problems aside for now and return to them in section 6.
@@ 5 The Relativistic Time Evolution Paradigm
Although it is not difficult to find a version of electromagnetism that fits the time evolution paradigm, that paradigm may be too loose. Electromagnetism is a relativistic theory and, as such, we can put tighter constraints on the form that the dynamical laws should take. They should be spatiotemporally local, [^26] Here I am talking about spatiotemporal locality in the special relativistic sense, not the broader sense in which there is no light-speed limit on influences and the only constraint is that there be spatiotemporal continuity (see Lange,, 2002, ch. 1). not merely temporally local. Here is how Maudlin, (2007, pg. 60) puts the requirement of relativistic locality: “the physical state at any point of space-time is determined or influenced only by events in its past light-cone.” [^27] See also Maudlin, (2007, pg. 20 & 30). If the state at any point can only be influenced by events in its past light-cone, then the state at a particular space-time point can only influence events in the future light-cone of that point (for these are the events that it is in the past light-cone of). Given Bell’s theorem, quantum physics may force us to give up on this kind of locality (a point that we will return to in section 7). But, it seems like this kind of locality should at least be achievable within classical electromagnetism.
@@
    Spatiotemporally Local Dynamical Laws: A subset of the laws (the dynamical laws) can be applied at any point in space-time, take as input only space-time point properties or past light-cone neighborhood properties, and give as output only past-to-future light-cone neighborhood properties. The dynamical laws may give either precise values for the past-to-future light-cone neighborhood properties or probability distributions over such values.
@@
    Locally Deterministic or Stochastic: Once a law-abiding history over an arbitrarily small time interval within the past light-cone of some future spacetime point has been fixed, the laws either uniquely fix a single future sequence of states within that light-cone or the laws yield a precise probability distribution over future sequences of states within that light-cone.
Present properties (that make no reference to other times) have been replaced by space-time point properties (that make no reference to other times or places). Past neighborhood properties have been replaced by past light-cone neighborhood properties (that are defined in terms of arbitrarily small subregions of the past light-cone). Past-to-future properties have been replaced by past-to-future light-cone properties. The third condition requires a local form of determinism or stochasticity. What is going on at a time within any spherical region (and its arbitrarily-short past light-cone) should either fix, or give probabilities for, what occurs within the remainder of the contracting light-cone that the spherical region is a slice of. [^28] For an example as to how such local determinism can be proven, see the analysis of the source-free homogeneous wave equation in Strauss, (2008, sec. 9.1).
@@ 6 Electromagnetism, Second Pass
It requires a bit of work to fit electromagnetism into the relativistic time evolution paradigm. The theory is normally understood as one where what happens at a given spacetime point is only influenced by its past light-cone and only influences its future light-cone, and thus normally understood to broadly fit the relativistic time evolution paradigm. However, precisely fitting the theory into that paradigm, as articulated in section 5, requires that the laws be spelled out with unusual care as to how the derivatives are understood so that, at every point in space-time, the laws link the past light-cone to the future light-cone. The situation here is similar to the case of Newtonian gravity (section 3), where the theory is normally understood as broadly fitting into the time evolution paradigm, but it takes some work to be careful about past and future derivatives so that the theory can be fit precisely into the time evolution paradigm from section 2. Electromagnetism is a more complex theory than Newtonian gravity, and the relativistic time evolution paradigm is a stricter standard than the time evolution paradigm, so more work is needed here than there.
Below I present one way to fit electromagnetism into the time evolution paradigm, without claiming that it is the best and final way. We need laws governing the evolution of the electromagnetic field, the forces on charged matter, and the reaction of matter to forces. Let us begin with the field.
6.1 The evolution of the electromagnetic field
We can use the scalar potential $\phi$ and the vector potential $\vec{A}$ to specify the state of the electromagnetic field, where these potentials are related to $\vec{E}$ and $\vec{B}$ by
@@ 6.1 The evolution of the electromagnetic field
Although this first attempt fails, we should be optimistic that the wave equations in (17) and (18) can be interpreted as giving spatiotemporally local dynamics. When the homogenous (source-free) wave equation, $\Box u=0$, is discussed in textbooks on partial differential equations, it is standard practice to prove a “causality theorem” showing that a given spacetime point cannot influence anything outside its future light-cone and cannot be influenced by anything outside its past light-cone. [^31] For discussion of causality theorems for one and three-dimensional wave equations, see Zachmanoglou & Thoe, (1976, ch. 8); Folland, (1995, ch. 5); Evans, (1998, sec. 2.4.3); Strauss, (2008, sec. 9.1). For discussion of causality theorems in electromagnetism (making use of the Lorenz gauge), see Wald, (1984, sec. 10.2); Wald, (2022, sec. 5.4). To be more precise, you can prove that if two solutions agree on the values of the function $u$ and its time derivatives $\frac{\partial u}{\partial t}$ at a given time within the spherical spatial region bounded by the past light-cone of some future spacetime point $(\vec{x},t)$, the two solutions will agree on the value of $u$ at $(\vec{x},t)$. Solutions to the inhomogeneous (sourced) wave equation, $\Box u=f$, will also satisfy such a causality theorem for a fixed source $f$ because any solution can be divided into a free part $u_{in}$ (that obeys the homogenous wave equation and can be interpreted as describing incoming waves that are not attributable to the source function $f$) and a retarded part $u_{ret}$ (where the value of $u_{ret}$ at any space-time point is fixed by $f$ along the past light-cone of that point). Given these causality results, it seems like there should be a way to formulate inhomogeneous wave equations, like the wave equations for the potentials (17) and (18), that makes the causal structure manifest and satisfies the conditions that we laid out earlier for spatiotemporally local dynamics linking past light-cone to future light-cone. As we will see, this can be done straightforwardly in one-dimensional space but is surprisingly difficult in three-dimensional space.
@@
with $f$ treated as a fixed source function defined across space and time. D’Alembert’s method for solving the homogeneous wave equation introduces new variables [^32] These are one-dimensional versions of Dirac’s light-cone, or light-front, coordinates (Dirac,, 1949).
@@
We can use these variables to rewrite (23) as
@@
where the derivatives are evaluated along the two perpendicular edges of the light-cone. If you were to take both derivatives to be future derivatives, then the output (the left-hand side of the equation) would be a future light-cone neighborhood property (and thus a past-to-future light-cone neighborhood property) and the input (the right-hand side) would be a space-time point property (assuming that $f$ gives space-time point properties). However, this interpretation of (25) would not capture the causal structure accurately. The source function $f$ alone does not determine how $u$ will change towards the future. At a moment, you need $f$, $u$, and the rate at which $u$ is changing to fix the future evolution.
@@
Looking back at (26), you could just as well swap the order of the derivatives and go up the other edge of the future light-cone. To get a more symmetric law and avoid privileging either side of the future light-cone, you could average the two orderings:
@@
where we will again treat $f$ as a fixed source function defined across space and time. The wave equations for the potentials, (17) and (18), fit this general form. Writing this wave equation as a spatiotemporally local dynamical law is not as straightforward as you might expect, but here is one way to do it. Let us focus our attention on the evolution at the origin, noting that any point in space can be treated as the origin. At the origin, [^34] Away from the origin, the Laplacian cannot be put in this form and must include derivatives with respect to $\theta$ and $\phi$ in addition to $r$ (as is evident in the standard expression for the Laplacian in spherical coordinates). the Laplacian $\nabla^{2}=\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}$ can be rewritten in spherical coordinates as an integral over contributions from second derivatives with respect to the radial coordinate $r$ for different choices of $\theta$ (the azimuthal angle) and $\phi$ (the polar angle)—seeing how $u$ changes as you look out in different directions from the origin: [^35] To derive (33), plug the expansion $\frac{\partial}{\partial r}=\cos\theta\sin\phi\frac{\partial}{\partial x}+\sin\theta\sin\phi\frac{\partial}{\partial y}+\cos\phi\frac{\partial}{\partial z}$ (31) into the integral over the solid angle. The cross terms drop out and you are left with $\int\frac{d\Omega}{4\pi}\left(3\frac{\partial^{2}}{\partial r^{2}}\right)u$ $=\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}\right)u=\nabla^{2}u\ .$ (32)
$\nabla^{2}u=3\int\frac{d\theta d\phi}{4\pi}\sin\phi\;\frac{\partial^{2}}{\partial r^{2}}u\ ,$ (33)
where $\int\frac{d\theta d\phi}{4\pi}\sin\phi$, or $\int\frac{d\Omega}{4\pi}$, integrates over the solid angle $\Omega$ and divides by the total solid angle. If we include a trivial integral over the solid angle for the time derivative term, the wave equation (30) for $u$ at the origin becomes
@@
which looks more like the one-dimensional wave equation (23).
@@
In the integral, for a given $\theta$ and $\phi$, the $\frac{\partial^{2}}{\partial\eta^{2}}$ term asks you to take a second derivative along the future light-cone (in the direction picked out by $\theta$ and $\phi$) and the $\frac{\partial^{2}}{\partial\xi^{2}}$ term asks you to take a second derivative along the past light-cone (in that direction). This past light-cone derivative should give the same result as taking the forward light-cone derivative in the opposite direction, $-\hat{r}$. Integrating over all angles, [^36] Here is a more formal proof of (38). Expanded in terms of $r$ and $t$ derivatives, the difference between the left and right-hand sides of (38) is proportional to $\int\frac{d\Omega}{4\pi}\left(\frac{\partial}{\partial r}\frac{\partial}{\partial t}\right)=0\ .$ (37) The fact that this term vanishes can be seen by noting that radial derivatives in opposite directions will cancel, or, by using the expansion of $\frac{\partial}{\partial r}$ in (31).
@@
Let us now distinguish past and future derivatives as in (26) and (29),
@@
For given angles $\theta$ and $\phi$, the integrand asks to consider how a sum of two past light-cone derivatives (one in the $\hat{r}$ direction and one in the $-\hat{r}$ direction) changes as you move along the future light-cone in the $\hat{r}$ direction (see figure 6). With this form of the wave equation, we have arrived at a spatiotemporally local dynamical law where the left-hand side is a past-to-future light-cone neighborhood property and the source function on the right-hand side is either a space-time point property or a past light-cone property (as we are about to see). [^37] One should not read the speed of causal influence directly from the ability to put the wave equation in a form like (41). If you started with a wave equation like (30) with the speed of light of light doubled ($c\rightarrow 2c$), local determinism would not hold (relative to ordinary light-cones) but you could follow parallel reasoning to the derivation above to get an expression for the wave equation that looks similar to (41) and uses the same coordinates $\eta$ and $\xi$ from (35).